OpenStudy (anonymous):

(-3x^3)(-2x^2) express using exponents

OpenStudy (mysesshou):

Are you trying to simplify it ? Like multiply it though? Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

simpli fy

OpenStudy (mysesshou):

\[(-3x^3)(-2x^2)\] So, the -3 * -2 gives you +6 \(x^3*x^2\) means you add the exponents here, giving \(x^5\) \[(-3x^3)(-2x^2)=6x^5\] So, if they are multiplied like this, then the exponents add. But if you had an exponent to an exponent, like \((x^3)^2=x^6\) then the exponents multiply does that help any ? :)

OpenStudy (anonymous):

yes it does thanks

OpenStudy (mysesshou):

yw